OpenStudy (anonymous):

Convert the given unit. 24 ft to yd A. 3yd B. 6yd C. 8yd D. 72yd P.s Help me and I will give you a Medal and fan.

OpenStudy (kirbykirby):

1 yd = 3 ft. So you can say: \[ \frac{1~yd}{3~ft}=\frac{x~yd}{24~ft}\] \(x = \frac{24\ti}{}

OpenStudy (kirbykirby):

\[ x = \frac{24\times 1}{3}=8\]